what is the table of values for the following equation 12x + -2y = -34
Answer by kkasko(46)

what is the table of values for the following equation 12x + -2y = -34?
Step 1: Solve for y
12x+-2y=-34
-12x -12x
-2y=-12x-34
/-2 /-2 /-2
y=6x+17
step 2: Create a chart x|y
put possible x values such as 0,1,2,3,4,5,6... in for x for instance
y=6(0)+17
y=0+17
y=17
y=6(1)+17
y=6+17
y=23
y=6(2)+17
y=12+17
y=29
x|y
0|17
1|23
2|29
3|35
4|41
5|47
6|53